I need to find an equation of the line satisfying the given condtion
Through (4,2); m= -2/3

Answer by benni1013(165)

Use slope-intercept form: y=mx+b
Info. Known: Point (4,2) m=-2/3
Your point is your x and y in your equation, so your equation set up should look like this:
2=4(-2/3)+b==>2=-8/3+b==>b=14/3
Now just re plug in:
y=(-2/3)x+14/3